Calculate the concentrations (give your answers to 3.s.f) of the following bases at 298K

(a) NaOH with a pH of 13.00

(b) Ba(OH)2 with a pH of 12.20


3. Calculate the pH (give your answers to 2.d.p) of the following bases.

(a) A solution of 40 g of sodium hydroxide dissolved in 500 cm3 of water at 298 K.

(b)80gdm–3 KOH at 298K



1
Expert's answer
2020-11-26T05:09:03-0500

1)     a)

pH+ pOH=14

13+ pOH=14

pOH=1

pOH=-log[OH-]

1=-log[OH-]

[OH-]= antilog [-1]

0.100M

b)

pOH=14-12.20

=1.8

-1.8=-log[OH-]

[OH-]=antilog (-1.8)

0.01584M

Ba(OH)2(s)=Ba2+(aq) +2OH-(aq)

Concentration of Ba(OH)2 =0.0.01584/2

=0.00792M

2)     a)

500cm3=40g

1000cm3=?

(1000/500) ×40

80g/li

Molarity=(g/li)/RMM

            =80/40

            =2M

pH=14+log2

                   =14.30

b)

molarity=80/56

            =1.4286M 

pH=14+log1.4286

14.15